OpenStudy (anonymous):

What sum is equal to 1/x+5 + x^2/x-5

OpenStudy (anonymous):

@Nnesha any ideas?

Nnesha (nnesha):

\[\huge\rm \frac{ 1 }{ x+5 } + \frac{ x^2 }{ x-5 }\] find the common denominator

Nnesha (nnesha):

find common denominator multiply the numerator of first fraction by the denominator of 2nd fraction multiply the numerator of 2nd fraction by the denominator of first fraction here is an example \[\huge\rm \frac{ a } {\color{Red}{ b} }+\frac{ c }{\color{blue}{ d} } =\frac{ a\color{blue}{d}+c\color{Red}{b} }{ bd}\]

OpenStudy (anonymous):

So 1x + 3x/ x+5 x-5

Nnesha (nnesha):

hmm that looks wrong when the denominators are not the same to get the common denominator multiply denominators of both fractions just like in the example i gave you

Nnesha (nnesha):

make sense ? or do you need an example with numbers ?

OpenStudy (anonymous):

yes please. I know this is a simple question but ive been doing algebra since about 8am lol my mind is jumbled

Nnesha (nnesha):

alright i'll give you an example \[\huge\rm \frac{ 2 }{ 3 } +\frac{ 4 }{ 5 }\] denominators aren't the same so i should multiply them 3 times 5 = 15 so 15 is my common denominator \[\frac{ ??? + ?? }{ 15 }\] when we convert two fractions to one we should multiply the numerator of `1st` fraction with the denominator of `2nd` fraction and multiply the numerator of `2nd` fraction with the denominator of first fracton just like cross multiplication but we should keep the positive sign